OpenStudy (anonymous):

Given the function T(z) = z – 6, find T(–4). (Points : 4) –10 –2 2 10

OpenStudy (freckles):

have you replaced z with -4?

OpenStudy (freckles):

you are only asked to perform the following: -4-6

OpenStudy (anonymous):

-10 or -2 maybe, I know for sure it will keep a negative and as a matter of fact I believe its

OpenStudy (anonymous):

maybe -

OpenStudy (freckles):

-4-6 is the same as -1(4+6)

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

its -10

OpenStudy (anonymous):

for sure?

OpenStudy (anonymous):

Becuase if you notice 4x-4 = -16 - 6 = -10

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

it cant be lower than 4 so u cross out 2

OpenStudy (anonymous):

cuz u multiply

OpenStudy (anonymous):

and it stays -